OpenStudy (anonymous):

For f(x) = 4x + 7, find f(4)

OpenStudy (anonymous):

If x in the f is 4... then x on the other side is also for so what is... f(4) = 4(4) + 7 ?

OpenStudy (anonymous):

four**

OpenStudy (anonymous):

4=16+7 4=23 f=19

OpenStudy (anonymous):

f(4) = 16+7 f(4) = 23

OpenStudy (anonymous):

You're in functions now :)

OpenStudy (anonymous):

oh ok got it thanks

OpenStudy (anonymous):

I would but I never heard of that :/